# AL-QASIDA: Analyzing LLM Quality and Accuracy Systematically in Dialectal Arabic

### v1

@@ 6 Conclusion
Acknowledgments
We thank Arwa Alaya, Noha Shehata, Abdullah Omran, Brenda Malacara, Linda Yanes, Sarah Laouedj, Djillali Boutouili, Asmaa Oirgari, Amir Hussein, Jay Alammar, Amr Keleg, and Nizar Habash for their invaluable contributions to the completion of this work.
@@ Appendix A Supplemental Figures and Graphics
Supplemental and more complete experimental results can be found in Figures 9 and 10 Figure 10 visualizes MT scores as in Figure 5, but for all eight dialects evaluated. We will host complete detailed evaluation results on a server upon formal publication of this work.
@@
We show the sizes of LLMs we used, in number of parameters, in Table 9. It is notable that Llama-3.1 scored best in the monolingual task, since we used only the 8B-parameter model, compared to much larger models Command-R+ and GPT-4o.

### v2

@@ 6 Conclusion
Lastly, we acknowledge that while all other evaluation sets we used were intended for this type of NLP evaluation, the song lyrics corpus we employed may have originally been intended for other purposes. Because most of these data sets were already curated for our general purposes, we did not vet them extensively for offensive content or sensative material. Our release of the song lyrics will not include any metadata (such as artist names), though such data can be found in its original source. Finally, we acknowledge use of LLMs to assist software development in this project, particularly in creating the graphics displayed in this paper.
@@ Appendix A Supplemental Figures and Graphics
Supplemental and more complete experimental results can be found in Figures 10 and 11 Figure 11 visualizes MT scores as in Figure 5, but for all eight dialects evaluated. We will host complete detailed evaluation results on a server upon formal publication of this work.
@@
We show the sizes of LLMs we used, in number of parameters, in Table 9. It is notable that Llama-3.1 scored best in the monolingual task, since we used only the 8B-parameter model, compared to much larger models Command-R+ and GPT-4o. We did not train any models and only performed inference.